Tom Schriber, a director for personnel of Management Resources, Inc., is in the process of designing...

  1. Tom Schriber, a director for personnel of Management Resources, Inc., is in the process of designing a program that its customers can use in the job-finding process. Some of the activities include preparing resumes, writing letters, making appointments to see prospective employers, researching companies and industries, and so on. Some of the information on the activities is shown in the following table:

Days

Immediate

Predecessors

Activity

a

m

b

A

8

9

10

__

B

6

7

8

__

C

3

3

5

A

D

10

20

30

B

E

6

8

8

B

F

9

10

12

C,D

G

6

8

10

E,C,D

  1. Determine the expected time and variance for each activity.
  2. Construct a network for this problem.
  3. Determine ES, EF, LS, LF, and slack for each activity.
  4. Determine the critical path and project completion time.
  5. Determine the probability that the project will be finished in 40 days or less. (15 points)
Answer #1

(a)

Activity a m b te = (a + 4m + b)/6 s2 = (b - a)2/36
A 8 9 10 9 0.111
B 6 7 8 7 0.111
C 3 3 5 3.333 0.111
D 10 20 30 20 11.111
E 6 8 8 7.667 0.111
F 9 10 12 10.167 0.250
G 6 8 10 8 0.444

(b)

A F Start Finish B D G E

(c)

Forward Pass:

ES of the starting activities = 0
ES of all other activities = Max. (EF of their immediate predecessors)
EF of an activity = Its ES + Its 'te'

Backward Pass:

LF of ending activities = Max. (All the EFs)
LF of all other activities = Min. (LS of their immediate successors)
LS of an activity = Its LF - Its 'te'

Total slack (or, float) = LF - EF for each activity.

Activity te = (a + 4m + b)/6 ES EF LS LF Slack
A 9 0.000 9.000 14.667 23.667 14.667
B 7 0.000 7.000 0.000 7.000 0
C 3.333 9.000 12.333 23.667 27.000 14.667
D 20 7.000 27.000 7.000 27.000 0
E 7.667 7.000 14.667 21.500 29.167 14.500
F 10.167 27.000 37.167 27.000 37.167 0
G 8 27.000 35.000 29.167 37.167 2.167

(d)

The critical path of a project is composed of activities with zero total slack.

So, the critical path is B-D-F. The expected completion time = 37.167 days.

(e)

Activity te = (a + 4m + b)/6 s2 = (b - a)2/36
B 7 0.111
D 20 11.111
F 10.167 0.250
Totals 37.167 11.472

Project's mean duration = 37.167 days

Stdev = sqrt(11.472) = 3.387 days

Target = 40 days

So, z = (40 - 37.167)/3.387 = 0.837

So, F(z) = NORMSDIST(0.837) = 0.799

So, the probability that the project is completed by 40 days = 0.799